OpenStudy (anonymous):

What is the standard form equation of the line passing through (5, –3) with a slope of –3?

OpenStudy (anonymous):

y=mx+b 5 is x -3 is y -3 is slope Just plug everything into the equation: -3=-3*5+b And then, find b. Finally rewrite the whole equation.

OpenStudy (anonymous):

Get it?

OpenStudy (anonymous):

Lol Wait Let Me Try Real Quick

OpenStudy (anonymous):

b= -18 ? .. . . . This Isnt Helpin Lol iNeed The Equation

OpenStudy (anonymous):

Ya, so you basically have everything for the equation now. Just rewrite the equation

OpenStudy (anonymous):

Wait, b doesn't equal -18 though. How'd you solve it?

OpenStudy (anonymous):

hi nitaleung is saying y=mx+b x=5 y=-3 slope m=-3 Just plug everything into the equation y=mx+b -3=-3*5+b solve for b b=15-3=12 so put it together in the formula y=mx+b y=-3x+12 ans